OpenStudy (mary.rojas):

What are all of the elements necessary to determine the equation of a line? writing a paper, any ideas?

OpenStudy (anonymous):

To determine the equation of a line, the two things you must need to find are: 1) slope of line 2) Any one point on the line Note: if the slope hasn't been given to you, you need at least 2 points along the line to find the slope.

OpenStudy (mary.rojas):

so talk about how i can find slope by looking at a graph, the b (y=mx+b), y intercepts, x intercepts, and idk what else

OpenStudy (anonymous):

Do you have a graph of the line?

OpenStudy (anonymous):

Do you know two points along the line? For example (3,1) and (-8,10) ? Have any been given?

OpenStudy (mary.rojas):

see thats what i dont get. our question is imply that:What are all of the elements necessary to determine the equation of a line? I guess it is up to us on how far we go. We can add a picture so I think I will add a simple graphed line

OpenStudy (anonymous):

Well, obviously you cannot determine an equation of the line if you only have one point. So, how i would explain it is by saying that you need any two points of the line. Lets say (0,0) and (2,6). To find the slope you use to formula y2-y1 / x2 - x1 = 6/2 = 3 Therefore, m=3 So in y=mx+c y = 3x + c To find c, simply substitute a point in the graph: 0=3(0) + c therefore c = 0 therefore equation is y=3x

OpenStudy (mary.rojas):

ahhhhh got you! Thank you, I will go with that route.

OpenStudy (anonymous):

Another way you can find the equation of the line is with the formula y - y1 = m (x - x1) So here, just substitute the slope as m and the point (2,6) into (x1, y1) and you will get the same answer. I like this method as it is quicker.

OpenStudy (anonymous):

@Thatyoungbuttercup